# Finding Most Shattering Minimum Vertex Cuts of Polylogarithmic Size in Near-Linear Time

### v1

@@ 9 Most Shattering Minimum Vertex Cut
We have presented a randomized near-linear time algorithm that lists all $k$-shredders with high probability, but we have not yet shown how to count the number of components each $k$-shredder separates. Counting components proves to be a bit trickier than it was in [CT99]. One may hope for the existence of an oracle that takes a set $S$ of $k$ vertices as input and counts the number of components in $G\setminus S$ in $\mathcal{O}(\text{poly}(k\log n))$ time. If there exists such an oracle, we are immediately finished because there are at most $n$ $k$-shredders. We can simply query such an oracle for each $k$-shredder in a total of $\mathcal{O}(n\text{poly}(k\log n))$ time. Unfortunately, such data structure cannot exist assuming SETH, as shown in [LS22, Theorem 8.9].
What we will do is modify our algorithms such that whenever we list a $k$-shredder $S$, we also return the number of components of $G\setminus S$. We will do this for Algorithm˜1, Algorithm˜2, and Algorithm˜3. We do not need to modify Algorithm˜7 because it lists $k$-shredders only by calling $\texttt{Shredders}(\cdot,\cdot)$ as a subroutine, and $\texttt{Shredders}(\cdot,\cdot)$ already supports counting the number of components as proved in [CT99]. Note that if there are no $k$-shredders of $G$, we can simply return a $k$-separator of $G$. This can be done in $\tilde{\mathcal{O}}(m+k^{3}n)$ time as shown in [FNS+20].
@@ 9.1 Counting Components in the Local Algorithm
Suppose that Algorithm˜1 returns $(\mathcal{L},U)$ on input $(x,\nu,\Pi)$. There exists a modified version of Algorithm˜1 such that for each $k$-shredder $S\in\mathcal{L}$ that is captured by $(x,\nu,\Pi)$, the modified version also computes the number of components of $G\setminus S$. The modification requires $\mathcal{O}(k^{2}\nu)$ additional time, subsumed by the running time of Algorithm˜1.
@@
Suppose Algorithm˜1 identifies a candidate $k$-shredder $S$ on line 19. By line 16, all bridges of $\Pi$ attached to some vertex $s\in S$ must have been explored. The modification is as follows. We can keep a dictionary $M$ whose keys are candidate $k$-shredders (in $k$-tuple form) mapped to nonnegative integers initialized to zero. For each bridge $\Gamma$ of $\Pi$ attached to $s$, if $\Gamma$ is a component of $G\setminus\Pi$ such that $N(\Gamma)=S$, then we increment $M[S]$. We prove that if $S\in\mathcal{L}$ is a $k$-shredder captured by $(x,\nu,\Pi)$, then $M[S]+2$ is the number of components of $G\setminus S$.
Let $S\in\mathcal{L}$ be a $k$-shredder with partition $(C,\mathcal{R})$ captured by $(x,\nu,\Pi)$. Let $Q_{x}$ denote the component in $\mathcal{R}$ containing $x$. From Lemma˜5.2, we have that every component in $\mathcal{R}\setminus\{Q_{x}\}$ is a component of $G\setminus\Pi$. This means every component in $\mathcal{R}\setminus\{Q_{x}\}$ is a bridge of $\Pi$. Furthermore, if we fix a vertex $s\in S$, notice that all components in $\mathcal{R}\setminus\{Q_{x}\}$ are bridges of $\Pi$ attached to $s$. These components are special; each component of $R\setminus\{Q_{x}\}$ is a bridge $\Gamma$ of $\Pi$ that satisfies $N(\Gamma)=S$. What our modification does is to count the number of bridges $\Gamma$ attached to $s$ that satisfy $N(\Gamma)=S$. This means that after processing all bridges attached to $s$, we have $M[S]=|\mathcal{R}\setminus\{Q_{x}\}|$. There are only two components of $G\setminus S$ that were not counted: $Q_{x}$ and $C$. Hence, there are $M[S]+2$ components of $G\setminus S$.
The time complexity for this operation is straightforward. For each candidate $k$-shredder $S$, we spend $\mathcal{O}(k)$ time to hash and store it in $M$. For each bridge $\Gamma$ of $\Pi$, we need to check whether $N(\Gamma)$ is a key in the dictionary. This can be done by hashing $N(\Gamma)$ in $\mathcal{O}(k)$ time. As proved in Lemma˜5.4, there are at most $\mathcal{O}(\nu)$ candidate $k$-shredders, and at most $\mathcal{O}(k\nu)$ explored bridges. Hence, hashing all candidate $k$-shredders and bridges can be done in $\mathcal{O}(k)\cdot\mathcal{O}(\nu+k\nu)=\mathcal{O}(k^{2}\nu)$ time. We conclude that these modifications impose an additional time cost of $\mathcal{O}(k^{2}\nu)$, which is subsumed by the runtime of the original algorithm. ∎
@@ 9.2 Counting Components for Low-Degree Unverified Sets
If Algorithm˜2 returned true on input $(x,\nu,\Pi,U)$, then $U$ is a $k$-shredder with a vertex $u\in U$ such that $\deg(u)\leq\nu$. Notice that all components of $G\setminus U$ must contain a vertex that is adjacent to $u$. Immediately, we have that the number of components of $G\setminus U$ is upper bounded by $\nu$, as $\deg(u)\leq\nu$. However, looking at two arbitrary edges $(u,v_{1}),(u,v_{2})$ adjacent to $u$, it is unclear whether $v_{1}$ and $v_{2}$ are in the same component of $G\setminus U$. Although we can query whether $v_{1}$ and $v_{2}$ are connected in $G\setminus U$ in $\mathcal{O}(k)$ time, we cannot afford to make these queries for all pairs of vertices in $N(u)$. Such a procedure would require us to make $\mathcal{O}(\nu^{2})$ many queries, which is too costly.
@@
Suppose that Algorithm˜2 returns true on input $(x,\nu,\Pi,U)$. There exists a modified version of Algorithm˜2 that also computes the number of components of $G\setminus U$. The modification requires $O(k^{2}\nu)$ additional time, subsumed by the running time of Algorithm˜2.
@@
We can remodel the problem of counting the components of $G\setminus U$ as the following. Fix a vertex $u\in U$, and suppose that $N(u)=\{y_{1},\ldots,y_{|N(u)|}\}$. For each vertex $y_{i}$, we wish to determine whether any vertex in $\{y_{1},\ldots,y_{i-1}\}$ is connected to $y_{i}$. Let $q$ denote the number of vertices $y_{i}$ such that $y_{i}$ is not connected to any vertex in $\{y_{1},\ldots,y_{i-1}\}$. It is straightforward to see that $q$ is the number of components of $G\setminus U$. The idea is to use a DFS-tree $T$ and calls to the pairwise connectivity oracle to count $q$. Consider Algorithm˜6.
@@
1: $(x,\nu,\Pi,U)$ - a tuple that Algorithm˜2 returned true on
2: the number of components of $G\setminus U$
@@
3: $T\leftarrow$ a global DFS-tree initialized prior to calling this algorithm
@@
4: $\mathcal{I}\leftarrow$ representatives for internal components of $T\setminus U$ (from Lemma˜9.4)
5:
$u\leftarrow$ a vertex in $U$ with $\deg(u)\leq\nu$
6:
7: $\texttt{counter}\leftarrow 0$
@@
8: for each edge $(u,y)$ adjacent to $u$ such that $y\notin U$ do
9:   $r_{y}\leftarrow$ the root of the component of $T\setminus U$ containing $y$
10:   if $r_{y}$ is marked as seen then
11:    skip to next edge
12:
mark $r_{y}$ as seen
13:
14:   $\texttt{flag}\leftarrow\texttt{true}$
15:   for each vertex $v\in\mathcal{I}$ do
16:    if $y$ and $v$ are not connected in $G\setminus U$ then
17:      skip to next vertex
18:    if $v$ is marked as seen then
19:      $\texttt{flag}\leftarrow\texttt{false}$
20:      break
21:    else
22:      mark $v$ as seen
23:
24:   if flag is true then
25:    $\texttt{counter}\leftarrow\texttt{counter}+1$
26: return counter
@@
Now we prove the correctness of Algorithm˜6. Suppose we are processing an edge $(u,y)$ during the for loop on line 8. We wish to determine whether there exists an edge $(u,y^{\prime})$ processed in a previous iteration such that $y$ and $y^{\prime}$ are connected in $G\setminus U$. Assume that such an edge exists. Let $(u,y^{\prime})$ be the earliest edge processed in a previous iteration such that $y$ and $y^{\prime}$ are connected in $G\setminus U$. Let $C_{y},C_{y^{\prime}}$ denote the components containing $y,y^{\prime}$ in $T\setminus U$, respectively. Lemma˜9.3 implies that either $C_{y}=C_{y^{\prime}}$, or $y$ and $y^{\prime}$ are both connected to an internal component of $T\setminus U$. If $C_{y}=C_{y^{\prime}}$, then we would have marked the root of $C_{y}$ as seen when processing $(u,y^{\prime})$ on line 12. Otherwise, $y$ and $y^{\prime}$ are both connected to an internal component $C$ of $T\setminus U$. Let $v$ be the unique vertex in $\mathcal{I}\cap C$, which we know must exist by Lemma˜9.4 and line 4. During the for loop on line 15, we will see that $v$ was marked as seen on line 18. This is because when processing $(u,y^{\prime})$, we must have eventually marked $v$ as seen on line 22. Conversely, suppose that there does not exist a prior edge $(u,y^{\prime})$ such that $y$ and $y^{\prime}$ are connected in $G\setminus U$. Again, Lemma˜9.3 implies that $C_{y}$ should not be marked as seen and no internal component connected to $y$ in $G\setminus U$ should be marked as seen. Otherwise, it would contradict our assumption. Therefore, flag is true by line 25 and we correctly increment the component counter.
@@
Now we prove the time complexity of the algorithm. Line 4 runs in $\mathcal{O}(k^{2})$ time by Lemma˜9.4. The main for loop on line 8 runs for $\mathcal{O}(\nu)$ iterations. Line 9 can be implemented in $\mathcal{O}(k)$ time as seen in [Kos23]. The for loop on line 15 runs for $\mathcal{O}(k)$ iterations. For each iteration, we need to query whether $y$ and $v$ are connected in $G\setminus U$ or whether $v$ has been marked as seen. This can be done in $\mathcal{O}(k)$ by querying the pairwise connectivity oracle in ˜6.2. Hence, lines 16-22 can be implemented in $\mathcal{O}(k)$. This means lines 15-22 runs in $\mathcal{O}(k^{2})$ time. Tallying all the steps up, the entire modification runs in $\mathcal{O}(k^{2}+k+\nu(k+k^{2}))=\mathcal{O}(k^{2}\nu)$ time additional to the time takes by Lemma˜6.6. ∎
@@ 9.3 Counting Components for High-Degree Unverified Sets
Suppose that Algorithm˜3 returns a list $\mathcal{L}$ of $k$-shredders on input $\mathcal{U}$. There exists a modified version of Algorithm˜3 such that for each $k$-shredder $S\in\mathcal{L}$, the modified version also computes the number of components of $G\setminus S$. The modification requires $\mathcal{O}(m\log^{2}n)$ additional time, subsumed by the time complexity of Algorithm˜3.
@@
Recall that Algorithm˜3 uses sampling to list high-degree $k$-shredders. Fix a high-degree $k$-shredder $S$ with partition $(C,\mathcal{R})$. Recall from Section˜6.2 that the key observation is that $S$ forms a wall of high-degree vertices. We exploited this by sampling vertices in $\mathcal{R}$ and exploring regions of vertices with low degree. The idea is that $S$ will restrict the graph exploration within one component of $\mathcal{R}$, and we can recover that component. We can actually extend this idea a bit further to count the number of components in $\mathcal{R}$. Let $Q$ be an arbitrary component of $\mathcal{R}$. Let $\nu$ be the power of two satisfying $\frac{1}{2}\nu<\mathrm{vol}(Q)\leq\nu$. By Lemma˜6.7, we can sample $400\frac{m}{\nu}\log n$ edges to obtain an edge $(y,z)$ such that $y\in Q$ with high probability. Since $\mathrm{vol}(\mathcal{R})<\frac{m}{k}$, we have $\mathrm{vol}(Q)<\frac{m}{k}$. So, if we vary $\nu$ over all powers of two up to $\frac{m}{k}$ and repeat the sampling scheme, we will eventually obtain a vertex $y\in Q$ for all components $Q\in\mathcal{R}$ with high probability. Notice that we have not modified Algorithm˜3 yet, the probability analysis already holds for the algorithm as it currently stands.
@@
The modification is as follows. We can keep a dictionary $M$ whose keys are the unverified sets in $\mathcal{U}$ (in $k$-tuple form) mapped to sets of vertices initialized to empty sets. We can think of $M[U]$ as a set of vertices that are each in a distinct component of $G\setminus U$. The idea is to add vertex samples to $M[U]$ one by one, collecting representatives in distinct components of $G\setminus U$. We add a few steps after line 13. First, we check whether any vertices in $Q$ are in the set $M[U]$. If so, this means that the component of $G\setminus U$ containing $y$ was already explored during a previous sample. Otherwise, we append $y$ to $M[U]$. We claim that if $U$ is a high-degree $k$-shredder in $\mathcal{L}$ at the end of Algorithm˜3, then the number of components of $G\setminus U$ is $|M[U]|+1$. The +1+1 term is due to the fact that the large component $C$ of $G\setminus U$ will not be counted because $\mathrm{vol}(C)\gg\frac{m}{k}$.
@@
To see why this modification works, notice that every time we enter the if branch on line 13, we have found a component $Q$ of $G\setminus U$ because $N(Q)=U$. The set $M[U]$ stores a list of vertices each in distinct components of $G\setminus U$ that have been explored. Whenever we explore a component $Q$, we simply need to check whether any vertices in $Q$ are in $M[U]$. If not, then it means $Q$ has not been explored before. Therefore, we add $y$ to $M[U]$ as a representative of the component $Q$. We have already showed that at the end of the entire algorithm, we will have sampled a vertex $x\in Q$ for all components $Q\in\mathcal{R}$ with high probability. This means all components in $\mathcal{R}$ will contain one unique representative in $M[U]$. Hence, $|M[U]|=|\mathcal{R}|$, so the number of components of $G\setminus U$ is $|M[U]|+1$.
@@
As for time complexity, we can use hashing to retrieve the set $M[U]$ given $U=N(Q)$ in $\mathcal{O}(k)$ time, as each unverified set must be exactly $k$ vertices. We can check whether a vertex is in $M[U]$ in $\mathcal{O}(1)$ time. Adding new vertices to a hash set can be done in $\mathcal{O}(1)$. Checking whether any vertex in $Q$ is in $M[U]$ can be done in $\mathcal{O}(\mathrm{vol}(Q))$ time by iterating over all vertices in $Q$ and performing a lookup in $M[U]$. Furthermore, the amount of insertion operations into $M[U]$ over all unverified sets $U$ is upper bounded by the number of vertex samples we use, which is at most $\mathcal{O}(m\log n)$ (recall Algorithm˜3). Since we spend $\mathcal{O}(\mathrm{vol}(Q))=\mathcal{O}(\nu)$ additional time for each sample, the additional cost of this modification is $\mathcal{O}(m\log^{2}n)$. We conclude that the cost is subsumed by the running time of the original algorithm. ∎
@@
We can keep track of the maximum number of components separated by all $k$-shredders listed so far. Using Lemmas˜9.1, 9.5 and 9.6, we can combine all three modifications so that for every $k$-shredder we list, we also keep track of the one maximizing the number of separated components. The proof of time complexity is identical to that of Lemma˜8.4. The only difference is that Algorithm˜2 now runs in $\mathcal{O}(k^{4}\log n+k^{2}\nu)$. However, this does not change the overall time complexity of the algorithm. ∎

### v2

@@ 9 Most Shattering Minimum Vertex Cut
We have presented a randomized near-linear time algorithm that lists all $k$-shredders with high probability, but we have not yet shown how to count the number of components each $k$-shredder separates. Counting components proves to be a bit trickier than it was in [CT99]. One may hope for the existence of an oracle that takes a set $S$ of $k$ vertices as input and counts the number of components in $G\setminus S$ in $\tilde{\mathcal{O}}(\text{poly}(k))$ time. If there exists such an oracle, we are immediately finished because there are at most $n$ $k$-shredders. We can simply query such an oracle for each $k$-shredder to spend $\tilde{\mathcal{O}}(\text{poly}(k)\cdot n)$ time in total. Unfortunately, such a data structure cannot exist assuming SETH, as shown in [LS22, Theorem 8.9].
What we will do is modify our algorithms such that whenever we list a $k$-shredder $S$, we also return the number of components of $G\setminus S$. We will do this for Algorithm 1, Algorithm 2, and Algorithm 3. We do not need to modify Algorithm 7 because it lists $k$-shredders only by calling $\texttt{Shredders}(\cdot,\cdot)$ as a subroutine, and $\texttt{Shredders}(\cdot,\cdot)$ already supports counting the number of components as proved in [CT99]. Note that if there are no $k$-shredders of $G$, we can simply return a $k$-separator of $G$. This can be done in $\tilde{\mathcal{O}}(m+k^{3}n)$ time as shown in [FNS+20].
@@ 9.1 Counting Components in the Local Algorithm
Suppose that Algorithm 1 returns $(\mathcal{L},U)$ on input $(x,\nu,\Pi)$. There exists a modified version of Algorithm 1 such that for each $k$-shredder $S\in\mathcal{L}$ that is captured by $(x,\nu,\Pi)$, the modified version also computes the number of components of $G\setminus S$. The modification requires $\mathcal{O}(k^{2}\nu)$ additional time, subsumed by the running time of Algorithm 1.
@@
Suppose Algorithm 1 identifies a candidate $k$-shredder $S$ on line 13. By line 12, all bridges of $\Pi$ attached to some vertex $s\in S$ must have been explored. The modification is as follows. We can keep a dictionary $M$ whose keys are candidate $k$-shredders (in $k$-tuple form) mapped to nonnegative integers initialized to zero. For each bridge $\Gamma$ of $\Pi$ attached to $s$, if $\Gamma$ is a component of $G\setminus\Pi$ such that $N(\Gamma)=S$, then we increment $M[S]$. We prove that if $S\in\mathcal{L}$ is a $k$-shredder captured by $(x,\nu,\Pi)$, then $M[S]+2$ is the number of components of $G\setminus S$.
Let $S\in\mathcal{L}$ be a $k$-shredder with partition $(C,\mathcal{R})$ captured by $(x,\nu,\Pi)$. Let $Q_{x}$ denote the component in $\mathcal{R}$ containing $x$. From Lemma 5.2, we have that every component in $\mathcal{R}\setminus\{Q_{x}\}$ is a component of $G\setminus\Pi$. This means every component in $\mathcal{R}\setminus\{Q_{x}\}$ is a bridge of $\Pi$. Furthermore, if we fix a vertex $s\in S$, notice that all components in $\mathcal{R}\setminus\{Q_{x}\}$ are bridges of $\Pi$ attached to $s$. These components are special; each component of $R\setminus\{Q_{x}\}$ is a bridge $\Gamma$ of $\Pi$ that satisfies $N(\Gamma)=S$. What our modification does is to count the number of bridges $\Gamma$ attached to $s$ that satisfy $N(\Gamma)=S$. This means that after processing all bridges attached to $s$, we have $M[S]=|\mathcal{R}\setminus\{Q_{x}\}|$. There are only two components of $G\setminus S$ that were not counted: $Q_{x}$ and $C$. Hence, there are $M[S]+2$ components of $G\setminus S$.
The time complexity for this operation is straightforward. For each candidate $k$-shredder $S$, we spend $\mathcal{O}(k)$ time to hash and store it in $M$. For each bridge $\Gamma$ of $\Pi$, we need to check whether $N(\Gamma)$ is a key in the dictionary. This can be done by hashing $N(\Gamma)$ in $\mathcal{O}(k)$ time. As proved in Lemma 5.4, there are at most $\mathcal{O}(\nu)$ candidate $k$-shredders, and at most $\mathcal{O}(k\nu)$ explored bridges. Hence, hashing all candidate $k$-shredders and bridges can be done in $\mathcal{O}(k)\cdot\mathcal{O}(\nu+k\nu)=\mathcal{O}(k^{2}\nu)$ time. We conclude that these modifications impose an additional time cost of $\mathcal{O}(k^{2}\nu)$, which is subsumed by the runtime of the original algorithm. ∎
@@ 9.2 Counting Components for Low-Degree Unverified Sets
If Algorithm 2 returned true on input $(x,\nu,\Pi,U)$, then $U$ is a $k$-shredder with a vertex $u\in U$ such that $\deg(u)\leq\nu$. Notice that all components of $G\setminus U$ must contain a vertex that is adjacent to $u$. Immediately, we have that the number of components of $G\setminus U$ is upper bounded by $\nu$, as $\deg(u)\leq\nu$. However, looking at two arbitrary edges $(u,v_{1}),(u,v_{2})$ adjacent to $u$, it is unclear whether $v_{1}$ and $v_{2}$ are in the same component of $G\setminus U$. Although we can query whether $v_{1}$ and $v_{2}$ are connected in $G\setminus U$ in $\mathcal{O}(k)$ time, we cannot afford to make these queries for all pairs of vertices in $N(u)$. Such a procedure would require us to make $\mathcal{O}(\nu^{2})$ many queries, which is too costly.
@@
Suppose that Algorithm 2 returns true on input $(x,\nu,\Pi,U)$. There exists a modified version of Algorithm 2 that also computes the number of components of $G\setminus U$. The modification requires $O(k^{2}\nu)$ additional time, subsumed by the running time of Algorithm 2.
@@
We can remodel the problem of counting the components of $G\setminus U$ as the following. Fix a vertex $u\in U$, and suppose that $N(u)=\{y_{1},\ldots,y_{|N(u)|}\}$. For each vertex $y_{i}$, we wish to determine whether any vertex in $\{y_{1},\ldots,y_{i-1}\}$ is connected to $y_{i}$. Let $q$ denote the number of vertices $y_{i}$ such that $y_{i}$ is not connected to any vertex in $\{y_{1},\ldots,y_{i-1}\}$. It is straightforward to see that $q$ is the number of components of $G\setminus U$. The idea is to use a DFS-tree $T$ and calls to the pairwise connectivity oracle to count $q$. Consider Algorithm 6.
@@
[…]
@@
1:$T\leftarrow$ a global DFS-tree initialized prior to calling this algorithm
@@
$\mathcal{I}\leftarrow$ representatives for internal components of $T\setminus U$ (from Lemma 9.4)
3:$u\leftarrow$ a vertex in $U$ with $\deg(u)\leq\nu$
4:
@@
5:$\texttt{counter}\leftarrow 0$Foreach edge $(u,y)$ adjacent to $u$ such that $y\notin U$
[…]
8:mark $r_{y}$ as seen
9:
[…]
17:return counter
[…]
@@
Now we prove the correctness of Algorithm 6. Suppose we are processing an edge $(u,y)$ during the for loop on line 5. We wish to determine whether there exists an edge $(u,y^{\prime})$ processed in a previous iteration such that $y$ and $y^{\prime}$ are connected in $G\setminus U$. Assume that such an edge exists. Let $(u,y^{\prime})$ be the earliest edge processed in a previous iteration such that $y$ and $y^{\prime}$ are connected in $G\setminus U$. Let $C_{y},C_{y^{\prime}}$ denote the components containing $y,y^{\prime}$ in $T\setminus U$, respectively. Lemma 9.3 implies that either $C_{y}=C_{y^{\prime}}$, or $y$ and $y^{\prime}$ are both connected to an internal component of $T\setminus U$. If $C_{y}=C_{y^{\prime}}$, then we would have marked the root of $C_{y}$ as seen when processing $(u,y^{\prime})$ on line 8. Otherwise, $y$ and $y^{\prime}$ are both connected to an internal component $C$ of $T\setminus U$. Let $v$ be the unique vertex in $\mathcal{I}\cap C$, which we know must exist by Lemma 9.4 and line 2. During the for loop on line 10, we will see that $v$ was marked as seen on line 11. This is because when processing $(u,y^{\prime})$, we must have eventually marked $v$ as seen on line 14. Conversely, suppose that there does not exist a prior edge $(u,y^{\prime})$ such that $y$ and $y^{\prime}$ are connected in $G\setminus U$. Again, Lemma 9.3 implies that $C_{y}$ should not be marked as seen and no internal component connected to $y$ in $G\setminus U$ should be marked as seen. Otherwise, it would contradict our assumption. Therefore, flag is true by line 16 and we correctly increment the component counter.
@@
Now we prove the time complexity of the algorithm. Line 2 runs in $\mathcal{O}(k^{2})$ time by Lemma 9.4. The main for loop on line 5 runs for $\mathcal{O}(\nu)$ iterations. Line 6 can be implemented in $\mathcal{O}(k)$ time as seen in [Kos23]. The for loop on line 10 runs for $\mathcal{O}(k)$ iterations. For each iteration, we need to query whether $y$ and $v$ are connected in $G\setminus U$ or whether $v$ has been marked as seen. This can be done in $\mathcal{O}(k)$ by querying the pairwise connectivity oracle in 6.2. Hence, lines 10-14 can be implemented in $\mathcal{O}(k)$. This means lines 10-14 runs in $\mathcal{O}(k^{2})$ time. Tallying all the steps up, the entire modification runs in $\mathcal{O}(k^{2}+k+\nu(k+k^{2}))=\mathcal{O}(k^{2}\nu)$ time additional to the time takes by Lemma 6.6. ∎
@@ 9.3 Counting Components for High-Degree Unverified Sets
Suppose that Algorithm 3 returns a list $\mathcal{L}$ of $k$-shredders on input $\mathcal{U}$. There exists a modified version of Algorithm 3 such that for each $k$-shredder $S\in\mathcal{L}$, the modified version also computes the number of components of $G\setminus S$. The modification requires $\mathcal{O}(m\log^{2}n)$ additional time, subsumed by the time complexity of Algorithm 3.
@@
Recall that Algorithm 3 uses sampling to list high-degree $k$-shredders. Fix a high-degree $k$-shredder $S$ with partition $(C,\mathcal{R})$. Recall from Section 6.2 that the key observation is that $S$ forms a wall of high-degree vertices. We exploited this by sampling vertices in $\mathcal{R}$ and exploring regions of vertices with low degree. The idea is that $S$ will restrict the graph exploration within one component of $\mathcal{R}$, and we can recover that component. We can actually extend this idea a bit further to count the number of components in $\mathcal{R}$. Let $Q$ be an arbitrary component of $\mathcal{R}$. Let $\nu$ be the power of two satisfying $\frac{1}{2}\nu<\mathrm{vol}(Q)\leq\nu$. By Lemma 6.7, we can sample $400\frac{m}{\nu}\log n$ edges to obtain an edge $(y,z)$ such that $y\in Q$ with high probability. Since $\mathrm{vol}(\mathcal{R})<\frac{m}{k}$, we have $\mathrm{vol}(Q)<\frac{m}{k}$. So, if we vary $\nu$ over all powers of two up to $\frac{m}{k}$ and repeat the sampling scheme, we will eventually obtain a vertex $y\in Q$ for all components $Q\in\mathcal{R}$ with high probability. Notice that we have not modified Algorithm 3 yet, the probability analysis already holds for the algorithm as it currently stands.
@@
The modification is as follows. We can keep a dictionary $M$ whose keys are the unverified sets in $\mathcal{U}$ (in $k$-tuple form) mapped to sets of vertices initialized to empty sets. We can think of $M[U]$ as a set of vertices that are each in a distinct component of $G\setminus U$. The idea is to add vertex samples to $M[U]$ one by one, collecting representatives in distinct components of $G\setminus U$. We add a few steps after line 9. First, we check whether any vertices in $Q$ are in the set $M[U]$. If so, this means that the component of $G\setminus U$ containing $y$ was already explored during a previous sample. Otherwise, we append $y$ to $M[U]$. We claim that if $U$ is a high-degree $k$-shredder in $\mathcal{L}$ at the end of Algorithm 3, then the number of components of $G\setminus U$ is $|M[U]|+1$. The $+1$ term is due to the fact that the large component $C$ of $G\setminus U$ will not be counted because $\mathrm{vol}(C)\gg\frac{m}{k}$.
@@
To see why this modification works, notice that every time we enter the if branch on line 9, we have found a component $Q$ of $G\setminus U$ because $N(Q)=U$. The set $M[U]$ stores a list of vertices each in distinct components of $G\setminus U$ that have been explored. Whenever we explore a component $Q$, we simply need to check whether any vertices in $Q$ are in $M[U]$. If not, then it means $Q$ has not been explored before. Therefore, we add $y$ to $M[U]$ as a representative of the component $Q$. We have already showed that at the end of the entire algorithm, we will have sampled a vertex $x\in Q$ for all components $Q\in\mathcal{R}$ with high probability. This means all components in $\mathcal{R}$ will contain one unique representative in $M[U]$. Hence, $|M[U]|=|\mathcal{R}|$, so the number of components of $G\setminus U$ is $|M[U]|+1$.
@@
As for time complexity, we can use hashing to retrieve the set $M[U]$ given $U=N(Q)$ in $\mathcal{O}(k)$ time, as each unverified set must be exactly $k$ vertices. We can check whether a vertex is in $M[U]$ in $\mathcal{O}(1)$ time. Adding new vertices to a hash set can be done in $\mathcal{O}(1)$. Checking whether any vertex in $Q$ is in $M[U]$ can be done in $\mathcal{O}(\mathrm{vol}(Q))$ time by iterating over all vertices in $Q$ and performing a lookup in $M[U]$. Furthermore, the amount of insertion operations into $M[U]$ over all unverified sets $U$ is upper bounded by the number of vertex samples we use, which is at most $\mathcal{O}(m\log n)$ (recall Algorithm 3). Since we spend $\mathcal{O}(\mathrm{vol}(Q))=\mathcal{O}(\nu)$ additional time for each sample, the additional cost of this modification is $\mathcal{O}(m\log^{2}n)$. We conclude that the cost is subsumed by the running time of the original algorithm. ∎
@@
We can keep track of the maximum number of components separated by all $k$-shredders listed so far. Using Lemmas 9.1, 9.5 and 9.6, we can combine all three modifications so that for every $k$-shredder we list, we also keep track of the one maximizing the number of separated components. The proof of time complexity is identical to that of Lemma 8.4. The only difference is that Algorithm 2 now runs in $\mathcal{O}(k^{4}\log n+k^{2}\nu)$. However, this does not change the overall time complexity of the algorithm. ∎